OpenStudy (anonymous):

If y varies directly as x and y = 28 when x = 4, then find k, the constant of proportionality.

OpenStudy (anonymous):

Is the answer 7?

OpenStudy (anonymous):

y=kx 28=k4 k=7

OpenStudy (anonymous):

If y varies inversely as x and y = 10 when x = 5, then the constant of variation would be: 2? @alyisnumber1

OpenStudy (anonymous):

y=k/x 10=k/5 k=50

OpenStudy (anonymous):

Thank you!